OpenStudy (vera_ewing):

Simplify completely: (x^2+4x-45)/(x^2+10x+9x)

Nnesha (nnesha):

first you have to factor top and bottom part

OpenStudy (vera_ewing):

\[\frac{ x^2+4x-45 }{ x^2+10x+9 }\]

OpenStudy (anonymous):

+9x?

OpenStudy (anonymous):

Ah

OpenStudy (anonymous):

Do as Nnesha says, factor the top and bottom :)

OpenStudy (vera_ewing):

What are the restrictions on the variable?

OpenStudy (anonymous):

What do you get when you factor the numerator?

OpenStudy (anonymous):

Hint: What adds up to 4, and what two products give you - 45.

OpenStudy (anonymous):

x-1?

OpenStudy (vera_ewing):

\[\frac{ x-5 }{ x+1 }\]

OpenStudy (anonymous):

Good

OpenStudy (vera_ewing):

What are the restrictions on the variable?

OpenStudy (anonymous):

Looks like -9 and -1

OpenStudy (vera_ewing):

Thank you.

OpenStudy (dtan5457):

@iambatman why -9?

OpenStudy (anonymous):

You have to still consider \[\frac{ (x+9)(x-5) }{ (x+9)(x+1) }\]